### v1

@@ 2 Related Work
Despite the potentially serious consequences of election harassment and the interdependent dynamics that can lead to the unraveling of fundamental democratic institutions; to the best of our knowledge, most existing literature only tangentially engages with these potential consequences, the specific role of technological shocks, and frameworks to rebuild resilience. One of the closest papers, by Dal Bo et al. [17], discusses the bribing and harassment of public officials as a two-stage game in which a population of prospective officials with heterogeneous qualities decides to apply for public office or enter the private sector. In the second stage, the interaction of public officials with pressure groups determines the among of public good provided. The authors investigate the equilibrium effects of bribes and threats and the consequences of policies such as official immunity. Other research focuses on strategic behaviors and incentives within election systems: Leontiou et al. [32] illustrate bandwagons in large and especially close elections, Andonie and Kuzmics [6] propose a two-stage game for opinion polls followed by elections, and find that the unique equilibrium for competing parties relates to their differing sample sizes. In empirical work, Berman et al. [9] discuss the fairness of elections in Afghanistan and [68] investigate a model of elections with candidate honesty and present experimental results using task performance as a laboratory model of policy making. In other works, Apffelstaedt et al. [4] and Le Moglie and Turati [33] look at the effects of elections on changing social norms and the news coverage of negative behaviors such as corruption, which is related to, but distinct from, targeted harassment that straddles the online and offline worlds.
@@ 4 Recruitment and Infiltration
Although new recruitment can limit the effect of harassment and increase resilience, we can also imagine a nefarious angle to recruitment: Consider “rogue" agents infiltrating the election system to replace leaving officials at each time step $t$, motivated, for example, by anecdotes of election deniers occupying such positions of influence over local elections [37]. We model the effect of rogue agents that replace the departing proportion $p_{t}$ as an increase in total harassment $H$ proportional to vacancies $p_{t}$ that can be abused by rogue agents who are not themselves targets of harassment, for example, because they are seen as sympathetic to the harassment goals [28]. In this case, let us denote the total harassment in the community at time $t$ by $H_{t}=H_{0}+\lambda p_{t}$, where $\lambda\geq 0$ measures the increasing rate of harassment from introduction of rouge agents proportionally to $p_{t}$, leading to the following attrition dynamics:
@@ 8 Discussion and Conclusions
Beyond the elections context, harassment reduces market efficiency and disrupts our organizational ability to efficiently aggregate information and collective decision making; e.g., public health experts facing harassment during the COVID-19 pandemic were unable to effectively guide public opinion towards the best public health practices. Although protesting is a democratic right, harassment crosses lines – especially if motivated by false rumors – that can uselessly disrupt the wisdom of crowds and waste scarce social capital. Building democratic resilience against harassment, while protecting the right to protest, is essential to the survival of critical civic infrastructure in the new (mis)information age.
@@ Appendix C A Stochastic Ordering
For general function $F:\mathbb{R}\rightarrow\boldsymbol{D\subset\mathbb{R}}$ , we know that if $F$ is continuous and injective, then it must be monotone(either a monotone increasing or a monotone decreasing function). If $F_{1}$ and $F_{2}$ are increasing functions, we know that $F_{1}(x)\leq F_{2}(x)$. Because $F_{1}^{-1}$ is increasing, $F_{1}^{-1}(F_{1}(x))\leq F_{1}^{-1}(F_{2}(x))$ and since $F_{1}^{-1}(F_{1}(x))=x=F_{2}^{-1}(g(x))$, we get $F_{2}^{-1}(F_{2}(x))\leq F_{1}^{-1}(F_{2}(x))$ or $F_{2}^{-1}(y)\leq F_{1}^{-1}(y)$, where $y=F_{2}(x)$, the inequality holds for all $y$. We know that $F_{1}$ dominating $F_{2}$ means that $F_{1}(x)\leq F_{2}(x)$ for all $x$, with a strict inequality at some $x$. Hence $F^{-1}_{2}\leq F^{-1}_{1}$, and we can compare their maximums if they exist. That is, from the above equation we know that $H_{1}=(1-p)Q_{1}(p)=(1-p)F^{-1}_{1}(p)\geq H_{2}=(1-p)Q_{2}(p)=(1-p)F^{-1}_{2}(p)$, that is, $H_{1}^{*}\geq H_{2}^{*}$. ∎

### v2

@@ 2 Related Work
Despite the potentially serious consequences of election harassment and the interdependent dynamics that can lead to the unraveling of fundamental democratic institutions; to the best of our knowledge, most existing literature only tangentially engages with these potential consequences, the specific role of technological shocks, and frameworks to rebuild resilience. One of the closest papers, by Dal Bo et al. [17], discusses the bribing and harassment of public officials as a two-stage game in which a population of prospective officials with heterogeneous qualities decides to apply for public office or enter the private sector. In the second stage, the interaction of public officials with pressure groups determines the amount of public good provided. The authors investigate the equilibrium effects of bribes and threats and the consequences of policies such as official immunity. Other research focuses on strategic behaviors and incentives within election systems: Leontiou et al. [32] illustrate bandwagons in large and especially close elections, Andonie and Kuzmics [6] propose a two-stage game for opinion polls followed by elections, and find that the unique equilibrium for competing parties relates to their differing sample sizes. In empirical work, Berman et al. [9] discuss the fairness of elections in Afghanistan and [68] investigate a model of elections with candidate honesty and present experimental results using task performance as a laboratory model of policy making. In other works, Apffelstaedt et al. [4] and Le Moglie and Turati [33] look at the effects of elections on changing social norms and the news coverage of negative behaviors such as corruption, which is related to, but distinct from, targeted harassment that straddles the online and offline worlds.
@@ 4 Recruitment and Infiltration
Although new recruitment can limit the effect of harassment and increase resilience, we can also imagine a nefarious angle to recruitment: Consider “rogue” agents infiltrating the election system to replace leaving officials at each time step $t$, motivated, for example, by anecdotes of election deniers occupying such positions of influence over local elections [37]. We model the effect of rogue agents that replace the departing proportion $p_{t}$ as an increase in total harassment $H$ proportional to vacancies $p_{t}$ that can be abused by rogue agents who are not themselves targets of harassment, for example, because they are seen as sympathetic to the harassment goals [28]. In this case, let us denote the total harassment in the community at time $t$ by $H_{t}=H_{0}+\lambda p_{t}$, where $\lambda\geq 0$ measures the increasing rate of harassment from introduction of rouge agents proportionally to $p_{t}$, leading to the following attrition dynamics:
@@ 8 Discussion and Conclusions
Beyond the elections context, harassment reduces market efficiency and disrupts our organizational ability to efficiently aggregate information and make decisions collectively; e.g., public health experts facing harassment during the COVID-19 pandemic were unable to effectively guide public opinion toward the best public health practices. Although protesting is a democratic right, harassment crosses lines – especially if motivated by false rumors – that can uselessly disrupt the wisdom of crowds and waste scarce social capital. Building democratic resilience against harassment, while protecting the right to protest, is essential to the survival of critical civic infrastructure in the new (mis)information age.
@@ Appendix C A Stochastic Ordering
Note that $F_{1}$ and $F_{2}$ are both increasing functions, and we know that $F_{1}(x)\leq F_{2}(x)$ for all $x$. Because $F_{1}^{-1}$ is increasing, $F_{1}^{-1}(F_{1}(x))\leq F_{1}^{-1}(F_{2}(x))$ and since $F_{1}^{-1}(F_{1}(x))=x=F_{2}^{-1}(F_{2}(x))$, we get $F_{2}^{-1}(F_{2}(x))\leq F_{1}^{-1}(F_{2}(x))$ or $F_{2}^{-1}(y)\leq F_{1}^{-1}(y)$, where $y=F_{2}(x)$, the inequality holding for all $y$. We know that $F_{1}$ stochastically dominating $F_{2}$ means that $F_{1}(x)\leq F_{2}(x)$ for all $x$, with a strict inequality at some $x$, and by monotonicity we showed that $F^{-1}_{2}(y)\leq F^{-1}_{1}(y)$ for all $y$, hence we can compare their supremum. That is, from the preceding equation we know that $(1-p)Q_{1}(p)=(1-p)F^{-1}_{1}(p)\geq(1-p)Q_{2}(p)=(1-p)F^{-1}_{2}(p)$ for all $p$, hence, $\mathcal{R}(F_{1})=\sup_{p\in(0,1)}(1-p)Q_{1}(p)\geq\mathcal{R}(F_{2})=\sup_{p\in(0,1)}(1-p)Q_{2}(p)$. ∎